Given: x2y+5xy+4y=0 Find y(2) given that the 2 arbitrary constants are equal and each has a value of 1.
Answer #1

zzy +5ay+ 4y = 0 bet x=12 a log x=2 - 2012 de gene la table > RD = 2 2 2 = 242=0, (loy) =) 22 = D. (0,-1) ① =) (0, (8,-1) +